# Determinations of angular stiffness in rotational optical tweezers

### v1

@@ V.1 The Helium-Neon Measurement Beam
Figure 4: The influence of the measurement beam on the equilibrium orientation of the probe and its angular trap stiffness for several probe sizes. (A) The shift in equilibrium angle is attributed to varying the numerical aperture of the measurement beam, which has a relative power of 0.1 of the trapping beam. (B) The shift in equilibrium and (C) relative change in angular trap stiffens, corresponding to the relative power of the measurement beam (NA 1.2).
In our experimental system, an ancillary measurement beam is introduced to probe the angular dynamics of the trapped particle. As described in Section II, this is a low-powered and weakly focussed circularly polarised $633\text{\,}\mathrm{nm}$ Helium–Neon laser. Owing to its low intensity and reduced numerical aperture, the measurement beam is treated as a passive probe of angular position that does not perturb the rotational dynamics of the probe. In this section, we investigate the validity and limitations of this assumption.
The circularly polarised measurement beam exerts a constant optical torque on the probe, which causes a shift in the equilibrium orientation of the probe in the direction of the applied torque. This displaces the probe from the zero-torque point of the trapping beam, corresponding to a reduction in the angular stiffness. However, a sufficiently large change in equilibrium orientation, $\Delta\phi_{0}$, can push the probe beyond the linear regime of the restoring torque, where stiffness estimates become less reliable. The measurement beam also affects the probe’s axial trapping, which varies the angular stiffness as optical torque is maximised near the beam focus. Hence, the combined influence of rotational and translational effects relative to the measurement beam’s parameters must be properly understood when interpreting angular dynamics.
Fig. 4 examines the influence of the He-Ne beam’s numerical aperture and its power relative to the trapping beam on the angular dynamics of the probe. Panel A shows the change in the equilibrium displacement as a function of its effective numerical aperture. The non-monotonic behaviour is specific to the rotational regime, as it is due to the coupling between the spin-angular momentum of the focussed beam and the inhomogeneous internal structure of the birefringent probe. Hence, the convergence angle that gives the maximum increase to stiffness is size-dependent.
@@
To explore the limits of this interaction, we consider the case where the relative power of the measurement beam is progressively increased beyond typical experimental conditions. At low beam convergence angles, increasing the He-Ne beam’s power can destabilise the trap entirely, as the gradient forces from both beams become insufficient to counteract the scattering force from the measurement beam. In such cases, no spatial trap is formed, and the probe escapes confinement. However, when a trap is formed, the constant torque from the circularly polarised beam can dominate the angular dynamics. As the torque from the measurement beam becomes comparable to or exceeds the restoring torque of the primary trap, the system enters a state of rotation instead of alignment. Hence, rotational dynamics are governed by the competition between alignment torque from the trapping beam and the rotation torque from the measurement beam.
Fig. 5 illustrates this behaviour, presenting a heat map of the relative power required to destabilise the system as a function of a probe size and measurement beam convergence angle. The line delineates the threshold between linear trap destabilisation and angular confinement failure. As expected, the axial trap fails to form at low numerical apertures, where the scattering force from the measurement beam dominates the total gradient force.
In addition to a destabilisation of angular trapping, there is also a relative beam power which causes the probe to depart from the translational regime. The departure from the linear regime causes the potential to become asymmetrical and no longer harmonic. Since the torque profile is a sinusoidal function, this is equivalent to asking: when does $\sin(2\phi)\approx 2\phi$, which can be defined as a sufficiently large shift in $\phi_{0}$. For example, setting $\phi_{0}>\pi/12$ as the departure from the linear regime requires a relative power between $0.48$ to $0.55$ times the power required to destabilise the alignment. These powers are an order of magnitude larger than typical experiments.
The relative power required to prevent angular trapping or depart from the translational regime is inversely related to the probe size. It becomes difficult to induce rotation for smaller probes, which has significant consequences for rotational measurements of nano- and sub-micron particles. These particles have low scattering cross-sections, which can drastically reduce the signal-to-noise in optical detection schemes. However, the power and convergence angle of the measurement beam can be optimised to enhance detection sensitivity without disrupting their expected dynamics. As shown in Fig. 5, angular trapping is maintained for smaller probes as the measurement beam power begins to surpass that of the trapping beam (provided axial trapping is achieved). This suggests that, contrary to conventional constraints, the measurement beam power can be increased to improve signal quality while preserving linear rotational dynamics.
Figure 5: A heat map showing the He-Ne measurement beam power relative to the trapping IR beam that destroys the measurement. The dashed line separates the regions into cases where either the trap or alignment cannot be formed.
@@ V.2 Spheroidal Vaterite Probes
The morphology of the vaterite probes varies between individual microspheres during synthesis, both in size and geometry. The probes have been observed to be slightly spheroidal, with the long axis aligned along their optic axis. To investigate how deviations from spherical geometry influence the angular trap stiffness, spheroidal vaterite probes were simulated with ellipticity introduced along the axis of rotational symmetry. The probe radius was defined along the fixed axes, and the aspect ratio was the length of the variable axis relative to this radius.
@@
In summary, deviations from spherical geometry introduce shape-induced birefringence that modifies the angular trap stiffness. This effect is size-dependent, with smaller probes exhibiting greater sensitivity to changes in aspect ratio due to enhanced curvature effects relative to the beam waist.
@@ V.3 Hydrodynamic and Inertial contributions
The theoretical expressions used in stochastic analysis techniques are derived from the equation of motion (Eq. 5), in which inertial and hydrodynamic terms are typically neglected. While this is a valid approximation for most optical trapping experiments, these effects become increasingly relevant at higher frequencies and can introduce systematic errors if omitted. In the translational case, berg-sorensen_power_2004 argued that hydrodynamic corrections should always be included when fitting the position power spectrum when precision is important. It is therefore necessary to reassess the relative contributions of these effects in the rotational regime.
@@
where $f_{i}$ is the characteristic rate associated with the inertia dissipation, and $\gamma(f)$ is the frequency-dependent drag coefficient that accounts for hydrodynamic effects. This expression (Eq. 27) is valid for both translational and rotational motion in one dimension, assuming appropriate substitutions as outlined in Table 1 of watson_interrogating_2025. The frequency-dependent drag coefficients for rotational and translational motion are given by
@@
In this section, we compare analytical power spectra generated by three models: the Lorentzian model, the hydrodynamically corrected theory, and the inertia-corrected theory, which incorporates both inertial and hydrodynamic effects. Since $f_{i}>f_{v}$, the hydrodynamic contribution must be included whenever inertial effects are considered. These models are shown in Fig. 7A-B for the translational and rotational regimes, respectively. The inertia-corrected model provides the most accurate description of the probe’s dynamics and is therefore the reference for comparison. The relative error between the Lorentzian and the hydrodynamically-corrected models, with respect to the inertia-corrected model, is shown in Fig. 7C for both translational and rotational motion. This figure provides a practical guide for determining which model is required to achieve a desired level of precision.
Regardless of the desired precision, the hydrodynamic and inertial contributions in the rotational are significantly smaller than their translational counterparts across most frequencies. In the rotational case, the combined contribution remains below 0.1% for frequencies up to 5.5 $5\text{\,}\mathrm{kHz}.$ By contrast, in the translational regime, the hydrodynamic contributions exceed 1% at only $13\text{\,}\mathrm{Hz}$, supporting the argument that it should always be included when fitting translational power spectra. This stark difference highlights the reduced sensitivity of rotational motion to fluid-mediated effects, which makes the simplified models—such as the Lorentzian—more broadly applicable in rotational optical trapping.
@@ VI Discussion
The calibration of rotational optical tweezers requires independent consideration of angular dynamics. Rotational and translational degrees of freedom appear analogous, due to the linearised model of their dynamics near equilibrium. This is demonstrated by the excellent agreement in determining angular stiffness from passive calibration techniques. Beyond this regime, the physical interactions governing rotational and translational motion differ notably. This study investigates important factors that determine angular trap stiffness, including probe geometry, hydrodynamic effects, and contributions from the ancillary measurement beam used in angular detection.
Optical tweezers are an attractive tool for investigating the mechanical properties of microscopic systems, owing to their ability to confine probes with high spatial precision. This capability is particularly advantageous for confined environments, such as microstructures and intracellular compartments. The use of rotational geometry and nanoparticles as probes further enhances the versatility of trapping experiments.
@@
The geometry and morphology of the probe have unique effects on the interaction with the optical potential and its underlying dynamics. The dynamics of spheroidal probes depend on the combination of shape-induced and material birefringence, and are amplified for smaller probes. Conversely, hydrodynamic and inertial effects are reduced, allowing them to be ignored in models of rotational dynamics at typical detection sensitivities.
@@
The parameters of both the trapping and measurement beams can influence the dynamics of the probe and the quality of the measurement. We demonstrated that the influence of the ancillary beam on the angular dynamics is minimal even when the beam parameters surpass the conventional limitations and decreases with probe size. This is particularly important for nanoparticles, which produce weak signals in optical detection schemes and make measurements of their dynamics more challenging. Unlike in the translational case, where increasing optical power alters the trapping dynamics, increasing the power of an ancillary beam in rotational optical tweezers enhances detection sensitivity without compromising the expected angular dynamics.
@@
In summary, these findings highlight the importance of treating translational and rotational dynamics as independent features of optical tweezers. Although their behaviour may initially appear analogous due to the linearised description of their dynamics near equilibrium, variations in experimental parameters yield profoundly different outcomes. Our work demonstrates the need for tailored approaches for calibrating and quantifying rotational optical tweezers independently from standard optical tweezers. This will lead to more accurate and targeted use of rotational optical tweezers in studies of complex microsystems, particularly towards nano-scale environments.

### v2

@@ V.1 The Helium-Neon Measurement Beam
Figure 4: Numerical calculations of the detection beam’s effect on vaterite probes of varying size for increasing beam power relative to the trapping beam. The panels show: (A) the change in the axial equilibrium position; (B) the change in the azimuthal equilibrium position with the dashed line indicated $15^{\circ}$; (C) the reduction in trap stiffness as a ratio with and without the trapping beam; and (D) the asymmetry of the angular stiffness represented by the ratio of the stiffness calculated by splitting the torque profile into the left and right sides to $\phi_{0}$. The detection beam convergence angle matches the trapping beam (equivalent to NA 1.2). The legend in panel C indicates the probe radius.
@@
The circularly polarised measurement beam exerts an optical force and an optical torque on the probe, modifying its dynamics depending on the beam’s power relative to the trapping beam. The optical force causes a slight shift to the axial equilibrium ($\Delta z_{0}$), and as demonstrated in Fig. 4A does not scale with probe size because the axial trap is sensitive to internal resonances. Since optical torque is maximised when the probe is at the beam focus, changes in the axial equilibrium will vary the angular trap stiffness. The additional optical torque vertically shifts the torque profile which moves the azimuthal equilibrium by $\Delta\phi_{0}$, displacing the probe from the zero-torque point of the trapping beam. This is shown in Fig. 4B. Unlike the change axial equilibrium, the change azimuthal equilibrium scales closely with probe size. These curves would be consistently distributed if each probe was position at the beam focus.
The equilibrium changes causes a reduction in the angular trap stiffness, as the azimuthal equilibrium is no longer at the maximum gradient of the alignment torque profile, as remonstrated by Fig. 4C. This reduction reflects the physical change in stiffness that occurs in experiments and is not an error into determining the $\chi$.
Instead, the main concern with a large $\Delta\phi_{0}$ is that an asymmetry is introduced, which can affect the fitting accuracy of the passive calibration models. The asymmetry is displayed in Fig. 4D as the ratio of the angular stiffness determined from only the left ($\phi<\phi_{0}$) and right ($\phi>\phi_{0}$) sides of the azimuthal equilibrium in the torque profile. A difference in $\chi_{\mathrm{left}}$ and $\chi_{\mathrm{right}}$ results in different relaxation times $\tau_{\mathrm{opt}}$ corresponding to increased confidence intervals from fitting using MSD, ACF, or PSD calibration methods. With these analysis methods, we would observe two curves corresponding to the left and right angular trajectories if they could be separated and measured with infinite precision. In reality, these curves are blurred together from the stochastic noise in the system resulting in larger confidence intervals from fitting routines. However, as the asymmetry remains very small ($1\%$ for realistic relative powers), this will only result in a slight increase in the confidence intervals from the fit.
As the measurement beam is not required for trapping, we can modify its convergence angle to alters the spin angular momentum transfer to the probe. This can improve the angular detection sensitivity [45], but also changes the amount of optical torque it exerts. Fig. 5 demonstrates the corresponding reduction in angular stiffness for a range of probe sizes when the detection beam has a relative power of 0.4 times that of the trapping beam. This relative power was chosen as it is slightly greater than typical experiments. These numerical results indicate a size-dependent non-monotonic relationship, which is unsurprising as the inhomogeneous and anisotropic structure of the vaterite probe has an optimum convergence angle for maximum angular momentum transfer.
Figure 5: The change in trap stiffness as a function of the detection beam convergence angle, with a relative power of 0.4 that of the trapping beam (NA 1.2). The legend indicates probe radius.
Following the results from Fig. 4 and Fig. 5, the measurement beam remains a passive sensor with negligible impact on trap stability under conventional operating beam parameters. Despite being largely negligible, these figures showed that the detection beam’s influence on angular trapping is more pronounced for larger sized probes. This is caused by the size-dependent competition between the alignment torque generated by the trapping beam and the constant rotation torque from the detection beam.
@@
Fig. 6 investigates the extremes of the interaction as a function of probe size and detection beam convergence angle, presenting a heat map of the relative power required to destabilise the system. At low beam convergence angles, increasing the He-Ne beam’s power can destabilise the spatial trapping, as the combined gradient forces from the two beams become insufficient to counteract the scattering force from the weakly-focussed measurement beam. In these cases, there is no spatial trap and the probe escapes confinement. When a trap is formed, the constant torque from the circularly polarised beam can dominate the angular dynamics. This occurs when the torque from the measurement beam exceeds the restoring torque of the primary trap, causing the system to enter a state of rotation instead of alignment. The white contour delineates the threshold between spatial trap destabilisation and angular confinement failure.
Figure 6: Numerical calculation of the relative power of the measurement He-Ne beam that prevents spatial trapping or alignment (separated by the white contour) given vaterite probe size and He-Ne beam convergence angle. The colour scale is logarithmic. The black contours indicate the azimuthal equilibrium shift when the beams have an equivalent power.
Since the torque profile is sinusoidal, the validity of the linear model is equivalent to the small angle approximation, where $\sin(2\phi)\approx 2\phi$. Hence, departure from the linear regime is arbitrarily defined based on the required precision. The black contours in Fig. 6 indicate the shift in equilibrium orientation due to the ancillary beam when it is at equal power to the trapping beam. As Fig. 4C showed $\Delta\phi_{0}\propto$ relative power, these contours serve as a reference guide to estimate $\Delta\phi_{0}$ for a given relative power and known probe size and ancillary beam convergence angle.
Fig. 4-6 highlight the competition between the alignment and constant rotation torques due to the two beams, and, more importantly, demonstrates the size-dependent relationship. Fig. 6 indicates that the relative power of the detection beam must increasingly exceed that of the trapping beam in order to disrupt the alignment of smaller particles—it becomes difficult to induce rotation for smaller probes. This has significant consequences for rotational measurements of nano- and sub-micron particles.
For example, a key problem with trapping experiments of nanoparticles is their low scattering cross-sections, which can drastically reduce the signal-to-noise in optical detection schemes. In ROTs experiments, increasing the detection beam power will increase the signal but has historically been avoided to maintain its negligible influence on dynamics. However, these results demonstrate that the detection beam power can be increased significantly, especially for smaller probes, without disrupting the angular dynamics—providing spatial trapping is still achieved. This provides an important method to improve signal quality while preserving rotational dynamics within the linear regime.
@@ V.2 Spheroidal Vaterite Probes
During synthesis, the morphology of the vaterite probes varies between individual microspheres, both in size and geometry. The probes have been observed to be slightly spheroidal, with the long axis aligned along their optic axis. To investigate how deviations from spherical geometry influence the angular trap stiffness, spheroidal vaterite probes were simulated with ellipticity introduced along the axis of rotational symmetry. The probe radius was defined along the fixed axes, and the aspect ratio was the length of the variable axis relative to this radius.
@@ V.3 Hydrodynamic and inertial contributions
The theoretical expressions used in stochastic analysis techniques are derived from the equation of motion (Eq. 5), in which inertial and hydrodynamic terms are often ignored. While this is a valid approximation for most optical trapping experiments, these effects become increasingly relevant at higher frequencies and can introduce systematic errors if omitted. In the translational case, Berg-Sørensen and Flyvbjerg [4] argued that hydrodynamic corrections should always be included when fitting the position power spectrum and precision is important. It is therefore necessary to reassess the relative contributions of these effects in the rotational regime.
@@
where $f_{i}$ is the characteristic rate associated with the inertia dissipation, and $\gamma(f)$ is the frequency-dependent drag coefficient that accounts for hydrodynamic effects. This expression (Eq. 26) is valid for both translational and rotational motion in one dimension, assuming appropriate substitutions as outlined in Table 1 of Watson et al. [44]. The frequency-dependent drag coefficients for rotational and translational motion are given by
@@
We compare analytical power spectra generated by three models: the Lorentzian model, the hydrodynamically corrected theory, and the inertia-corrected theory, which incorporates both inertial and hydrodynamic effects. Since $f_{i}>f_{v}$, the hydrodynamic contribution must be included whenever inertial effects are considered. These models are shown in Fig. 8A-B for the translational and rotational regimes, respectively. The inertia-corrected model provides the most accurate description of the probe’s dynamics and serves as the reference for comparison. The relative error between the Lorentzian and the hydrodynamically-corrected models, with respect to the inertia-corrected model, is shown in Fig. 8C for both translational and rotational motion. This figure provides a practical guide for determining which model is required to achieve a desired level of precision.
The hydrodynamic and inertial contributions in the rotational regime are significantly smaller than their translational counterparts across most frequencies. In the rotational case, the combined contribution remains below 0.1% for frequencies up to 5.5 $5\text{\,}\mathrm{kHz}.$ By contrast, the translational hydrodynamic contributions exceed 1% at only $13\text{\,}\mathrm{Hz}$, supporting the established argument that it should always be included when fitting translational power spectra. This stark difference highlights the reduced sensitivity of rotational motion to fluid-mediated effects, which makes the simplified models—such as the Lorentzian—more broadly applicable in rotational optical trapping.
@@ VI Discussion
In the linear regime near equilibrium, both translational and rotational stiffness determinations rely on analogous models, enabling passive analysis methods to recover trap stiffness with excellent consistency. However, this equivalence holds only for determining trap stiffness with the linear model of the respective one-dimensional dynamics. When examining how experimental parameters influence system behaviour, the physical interactions governing rotational and translational motion differ notably. As this study demonstrates, factors such as probe geometry, hydrodynamic coupling, and the influence of the auxiliary detection beam shape the angular trap stiffness and rotational dynamics in ways that have no direct translational analogue.
These results expand the versatility of rotational trapping experiments and are particularly valuable for studies of mechanical properties in confined environments, such as microstructures and intracellular compartments. A vital aspect to reaching smaller systems is the use of nanoparticles as rotational probes, which can leverage the results demonstrated across Section V to gain useful functionality.
@@
The geometry and morphology of the probe have unique effects on the interaction with the optical potential and its underlying dynamics. The angular dynamics of spheroidal probes depend on the combination of shape-induced and material birefringence, and are amplified for smaller probes. Conversely, hydrodynamic and inertial effects are reduced, allowing them to be ignored in models of rotational dynamics at typical detection sensitivities.
@@
The parameters of both the trapping and measurement beams can influence the dynamics of the probe and the quality of the measurement. We demonstrated that the influence of the ancillary beam on the angular dynamics is minimal even when the beam parameters surpass the conventional limitations and decreases with probe size. This is significant for nanoparticles, which produce weak signals in optical detection schemes making measurements of their dynamics more challenging. Unlike in the translational case, where increasing optical power alters the trapping dynamics, we show that increasing the power of an ancillary beam in rotational optical tweezers does not compromise the alignment dynamics allowing for enhanced detection sensitivity.
When viewed in combination with the geometric, hydrodynamic, and inertial effects, rotational optical tweezers experiments with nanoparticles are becoming increasingly attractive. Progress towards these advantages is currently only limited by the optical probe, as many spherical birefringent probes become less practical on sub-micron scales.
@@
In summary, these findings highlight the importance of treating translational and rotational dynamics as independent features of optical tweezers. Although their behaviour may initially appear analogous due to the linearised description of their dynamics near equilibrium, variations in experimental parameters yield profoundly different outcomes. Our work demonstrates the need for tailored approaches quantifying rotational optical tweezers independently from standard optical tweezers. This will lead to more accurate and targeted use of rotational optical tweezers in studies of complex microsystems, particularly towards nano-scale environments.